Calculate the elasticity of supply if a 15% increase in the price of soya bean oil increases its supply from 300 to 345 units


1
Expert's answer
2021-04-13T07:08:01-0400

es = % change in supply / % change in price

change in supply = (100 x (Q1-Q)) / Q = (100 x (345-300)) / 300 = 15%

es = 15 / 15 = 1